OpenStudy (anonymous):

Divide. (-6m^7 + 20m^6) ÷ 2m^4 A. -3m^7 + 10m^6 B. -6m^7 + 20m^2 C. -3m^3 + 10m^2 D. -3m^3 + 20m^6

OpenStudy (anonymous):

\[\frac{-6m^7+20m^6}{2m^4}\] \[= \frac{2m^4(-3m^3+10m^2)}{2m^4}\] \[=-3m^3+10m^2\]

OpenStudy (anonymous):

\[(-6m^7 + 20m^6) * \frac{1}{2m^4} \]

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

No worries :)